OpenStudy (accidentalaichan):

Review? 3x − 6y = 3 7x − 5y = −11 I need to find x, but I'm having trouble with negative integers. 10x + -1y = -7?

jigglypuff314 (jigglypuff314):

you would make the y values equal by multiplying them through to get to a LCM 3x − 6y = 3 -> 15x - 30y = 15 7x − 5y = −11 -> 42x - 30y = -66 now can you try? :)

OpenStudy (accidentalaichan):

okay, the -30y and-30y doesn't cancel out? -66 + 15 = -51? 15x + 42 x = 57x 57x - 30 y = -51

jigglypuff314 (jigglypuff314):

mmm? 3x − 6y = 3 -> 15x - 30y = 15 7x − 5y = −11 -> 42x - 30y = -66 so 42x - 30y = -66 -15x + 30y = -15 that look better for you? :3

OpenStudy (accidentalaichan):

Hmm, y = -2 I need to combine these ;-; 10 x - 2 = -7?

OpenStudy (accidentalaichan):

It's really confusing to me

jigglypuff314 (jigglypuff314):

errr, what's your question?

OpenStudy (texaschic101):

is this 2 problems ?

OpenStudy (accidentalaichan):

It's a system of equations

OpenStudy (texaschic101):

with 3 equations ?

OpenStudy (accidentalaichan):

10x = -5 so x = 2?

OpenStudy (accidentalaichan):

No, it's 2 equations

OpenStudy (accidentalaichan):

*-2

jigglypuff314 (jigglypuff314):

ummm could we back track a little? 42x - 30y = -66 -15x + 30y = -15 what did you get for this?

OpenStudy (accidentalaichan):

27x = -81

jigglypuff314 (jigglypuff314):

yep :) then divide both sides by 27 to get x

OpenStudy (accidentalaichan):

x = -3?

jigglypuff314 (jigglypuff314):

yep! :) then plug in what you got for x back into one of the original equations to get y so plug x= -3 into 3x − 6y = 3

OpenStudy (accidentalaichan):

3(-3) - 6(-2) = 3 -9 - (-12) = 3?

jigglypuff314 (jigglypuff314):

mmm don't plug anything else in for y? 3(-3) - 6y = 3 solve for y

jigglypuff314 (jigglypuff314):

oh wait, did you already get y? whoops, then ignore me :P

OpenStudy (accidentalaichan):

-9 - 6y = 3 +9 <- cancels -6y = 12 y = -2?

OpenStudy (accidentalaichan):

You're such a life saver ;u; Kindness bless your heart~

jigglypuff314 (jigglypuff314):

glad I could help :) <3

OpenStudy (accidentalaichan):